what is the answer to y=x+2 & y=2x+7 this is systems of linear equations?
Answer by stanbon(75887)

what is the answer to y=x+2 & y=2x+7 this is systems of linear equations?
-----
Equation:
x+2 = 2x+7
x = -5
=====
Solve for "y":
y = x+2
y = -5+2 = -3
-------------------
Ans: (-5, -3)
================
Cheers,
Stan H.